OpenStudy (anonymous):

sue, an experienced shipping clerk, can fill a ertain order 9 hours. Jim, a new clerk, needs 10 hours to do the same job. working together how long will it take them to fill the order?

OpenStudy (anonymous):

\[\frac{ order }{ 9 hours } *x hours +\frac{ order }{ 10 hours } *xhours =1\] or \[\frac{ x }{ 9 } + \frac{ x }{ 10 } =1\]

OpenStudy (anonymous):

use method above, or go right to \[\frac{8\times 9}{8+9}\] as a solution

OpenStudy (anonymous):

?

OpenStudy (anonymous):

why 8 and 9?

OpenStudy (anonymous):

because it is late an i am tired

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[\frac{9\times 10}{9+10}\] might work better really if you do more than two of these style problems it is better to go right to the answer